# Computing in a Faulty Congested Clique

### v1

@@ 5 Application: Semi-Ring Matrix Multiplication
Figure 2: Illustration for the inner and outer partitions, demonstrated on matrix $A$.
@@ 6 Leveraging a Sublinear Bound on the Number of Faults
    In the subsequent $c=O(1)$ quiet rounds, each node partitions its input into $n^{1-\chi}$ parts, each of size $n^{\chi}$. For each part, it forms an MDS code with parameters $[n^{\chi},\frac{n^{\chi}}{c},\frac{c-1}{c}n^{\chi}+1]$. It sends one codeword to each of $n^{1-\chi}$ subsets of size $n^{\chi}$ of a predetermined partition of the nodes which they all use.
@@
    The recursive steps remain almost the same, with the exception that the codewords now have parameters $[n^{\chi},\frac{n^{\chi}}{c},\frac{c-1}{c}n^{\chi}+1]$ and that the $n$ nodes do not hold one codeword but instead are partitioned into holding $n^{1-\chi}$ disjoint codewords of length $n^{\chi}$. In particular, the only thing that changes is that $(\widetilde{P}_{\ell,v}^{(k)})_{u}$ is now of size $n^{\chi}$ and $u$ also takes the more restricted range of $n^{\chi}$, since this is the length of the codewords.
The proof of correctness remains the same up to accordingly replacing $n$ with $n^{\chi}$ in appropriate places. E.g., for the the inductive hypothesis, we assume that epoch $\ell$ has been correctly checkpointed as at most $O(n^{\zeta+1-\chi})$ codewords of an $[n^{\chi},\frac{n^{\chi}}{c},\frac{c-1}{c}n^{\chi}+1]_{q}$-code for each node. In particular, although there are $O(n^{\zeta+1-\chi})$ codewords, they can be collected concurrently in “batches” of size $n^{1-\chi}$ and, likewise, collecting the data from any of these “batches” cannot fail since the adversary is constrained to fail at most $\frac{c-1}{c}n^{\chi}$ nodes. ∎
@@ 7 Application: Fast (Ring) Matrix Multiplication
Suppose that the inputs on the nodes are matrices $A,B\in\Sigma^{n\times n}$, where each node begins with $n$ coefficients of $A$ and $B$ and $\Sigma$ is a ring, then the value $C=A\cdot B$ can be computed in the ($\chi,c$)-$\mathsf{Faulty~{}Clique~{}}$with $O(1)$ quiet rounds, a round complexity of $O(\max\{c^{2}n^{\chi}\log n,c^{2}n^{1-2/\omega}\log n\})$, and $O(1)$-decodability, where $\omega$ is the exponent of matrix multiplication.
@@
$\begin{split}\widehat{A}_{k}&=\sum_{i,j\in[n^{1/\sigma}]}\alpha^{k}_{i,j}A^{i}_{j}\\ \widehat{B}_{k}&=\sum_{i,j\in[n^{1/\sigma}]}\beta^{k}_{i,j}B^{i}_{j}.\\ C^{i}_{j}&=\sum_{k\in[n]}\gamma^{i,j}_{k}\widehat{A}_{k}\widehat{B}_{k}.\\ \end{split}$ (8)
@@
$\begin{split}C^{i}_{j}[m,p]&=\sum_{k\in[n]}\gamma^{i,j}_{k}\left(\widehat{A}_{k}\widehat{B}_{k}\right)[m,p].\\ \end{split}$ (9)

### v2

@@ 5 Application: Semi-Ring Matrix Multiplication
Figure 2: Illustration for the inner and outer partitions, demonstrated on matrix AA.
@@ 6 Leveraging a Sublinear Bound on the Number of Faults
    In the subsequent $c=O(1)$ quiet rounds, each node partitions its input into $n^{1-\chi}$ parts, each of size $n^{\chi}$. For each part, it forms an MDS code with parameters $n^{\chi},\frac{n^{\chi}}{c},\frac{c-1}{c}n^{\chi}+1]$. It sends one codeword to each of $n^{1-\chi}$ subsets of size $n^{\chi}$ of a predetermined partition of the nodes which they all use.
@@
    The recursive steps remain almost the same, with the exception that the codewords now have parameters $n^{\chi},\frac{n^{\chi}}{c},\frac{c-1}{c}n^{\chi}+1]$ and that the $n$ nodes do not hold one codeword but instead are partitioned into holding $n^{1-\chi}$ disjoint codewords of length $n^{\chi}$. In particular, the only thing that changes is that $(\widetilde{P}_{\ell,v}^{(k)})_{u}$ is now of size $n^{\chi}$ and $u$ also takes the more restricted range of $n^{\chi}$, since this is the length of the codewords.
The proof of correctness remains the same up to accordingly replacing $n$ with $n^{\chi}$ in appropriate places. E.g., for the the inductive hypothesis, we assume that epoch $\ell$ has been correctly checkpointed as at most $O(n^{\zeta+1-\chi})$ codewords of an $n^{\chi},\frac{n^{\chi}}{c},\frac{c-1}{c}n^{\chi}+1]_{q}-$code for each node. In particular, although there are $O(n^{\zeta+1-\chi})$ codewords, they can be collected concurrently in “batches” of size $n^{1-\chi}$ and, likewise, collecting the data from any of these “batches” cannot fail since the adversary is constrained to fail at most $\frac{c-1}{c}n^{\chi}$ nodes. ∎
@@ 7 Application: Fast (Ring) Matrix Multiplication
See 3
@@
[…]
@@
[…]